OpenStudy (anonymous):

How far can you pull a wagon using 5N of force doing 50 J of work? A: 55m B: 250 m C: 45m D: 10m

OpenStudy (jfraser):

Since work is equal to force * distance, you're given both work and force, so solve for distance\[W = F*d\]so\[d = \frac{W}{F}\]